OpenStudy (anonymous):

Julio invests $500 each quarter in a fixed-interest mutual fund paying annual interest of 5% compounded quarterly. How much will her account have in it at the end of 15 years?

OpenStudy (anonymous):

\[500(1+\frac{.05}{4})^{15\times 4}\]

OpenStudy (anonymous):

then a calculator for sure

OpenStudy (anonymous):

i was using that one but i dont get it

OpenStudy (anonymous):

don't get it as in don't get the right answer or don't get it as in don't get where the equation comes from?

OpenStudy (anonymous):

Answers: $31,347.99 $41,348.01 $44,287.25 $51,055.59

OpenStudy (anonymous):

this is what i get http://www.wolframalpha.com/input/?i=500%281%2B.05%2F4%29^%2860%29

OpenStudy (anonymous):

oh i should learn how to read. it says invests 500 EACH QUARTER

OpenStudy (anonymous):

yeah that's my answer on my ti too

OpenStudy (anonymous):

scratch what i wrote

OpenStudy (anonymous):

\[A=P*((1+i)^n-1)/i\]

OpenStudy (anonymous):

sorry i mis interpreted the problem he invests 500 per quarter so again you have so sum a geometric series 500 \(500+500(.1025)\) \(500+500(.1025)+500(1.025)^2\) etc

OpenStudy (anonymous):

so i guess it is \[500(1+1.0125+(1.0125)^2+...\] is that rigth?

OpenStudy (anonymous):

\[\frac{1.0125^{59}-1}{1.0125-1}\] \[\frac{1.0125^{59}-1}{.0125}\]

OpenStudy (anonymous):

oh yeah, it is what you wrote, i just rederived it. ok lets computer \[500\frac{1.0125^{60}-1}{0.0125}\]

OpenStudy (anonymous):

looks like we got a winner. 44287.

OpenStudy (anonymous):

sorry i didn't read correctly at the beginning always helps to read

OpenStudy (anonymous):

yeah correct bro i appreciate it thank u

OpenStudy (anonymous):

yw